nine is 3/5 of what number? 3,15,27,or none of these? thanks for the help!!!!!!!!!!!!!!!!!!!!!!!
Answer by sabanasir(22)

3/5 of _______ = 9
suppose the empty space as 'x' meaning unknown.
3/5 of x = 9
now multiply on both sides the multiplicative inverse of 3/5 i.e. 5/3
3/5 * 5/3 of x = 9 * 5/3
x = 3 * 5
x = 15
hope u got it.